OpenStudy (anonymous):

fan and medal #PleaseeeHelp!!! @ganeshie8 pleas help me. An expression is shown below: 3x3y + 12xy − 9x2y − 36y Part A: Rewrite the expression so that the GCF is factored completely. Show the steps of your work. (4 points) Part B: Rewrite the expression completely factored. Show the steps of your work. (6 points)

OpenStudy (anonymous):

I got you

OpenStudy (anonymous):

So do you know what the GCF is?

OpenStudy (anonymous):

@EmilyF03

OpenStudy (anonymous):

greatest common factor, I think

OpenStudy (anonymous):

And how do you find?

OpenStudy (anonymous):

For this question the GCF is 3, do you know why? I am sorry I have to ask questions like this. I want you to think on your own, not just give you the answer.

OpenStudy (anonymous):

@EmilyF03 you need to participate...

OpenStudy (anonymous):

Gtg sorry

OpenStudy (anonymous):

sorry my comp is so slow... can with continue please @twistnflip

OpenStudy (anonymous):

are there options? multiple choice?

OpenStudy (anonymous):

no there isn't multiple choices

OpenStudy (anonymous):

First combine like terms. What are the like terms? List the like terms

OpenStudy (anonymous):

Please do not message me. It is ok if you forgot how to find like terms

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Like terms are terms that are a like. For instance, 3z3y + 4z6y are like terms. 4a5b + 7b are not like terms. What are the like terms?

OpenStudy (anonymous):

Like terms are terms that are a like. For instance, 3z3y + 4z6y are like terms. 4a5b + 7b are not like terms. What are the like terms?

OpenStudy (anonymous):

if I combine these like terms 3z3y + 4z6y I would get 3z3y + 4z6y = 7z9y

OpenStudy (anonymous):

so what I need to do now to find the answer for part a? or I need to cobine like terms now?

OpenStudy (anonymous):

@nixy

OpenStudy (anonymous):

\[3y \left( x^3-3x^2+4x-12 \right)\] \[=3y \left\{ x^2\left( x-3 \right)+4\left( x-3 \right) \right\}\] \[=3y \left( x-3 \right)\left( x^2+4 \right)\]

OpenStudy (anonymous):

its ht the answer for pat A? @surjithayer

OpenStudy (anonymous):

i don't know the answer GCF=3y

OpenStudy (anonymous):

what about first line for A?

OpenStudy (anonymous):

i have not rewritten first step of rearranging only thing is i have taken GCF out